Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability questioon

Expert replies
by melon » Sat Jan 28, 2012 12:05 am
A bag contains 3 white balls, 3 black balls & 2 red balls. One by one three balls are drawn out without replacement. What is the probability that the third ball is red?
Join the discussion
Source: — Problem Solving |

by LalaB » Sat Jan 28, 2012 2:21 am
since any ball has equal chances to be white, back or red, then the probability that the third ball is red is 2/8=1/4 (please note that 8 is the sum of all balls , i.e. 3 white balls, 3 black balls & 2 red balls 3+3+2=8)
Join the discussion

by Anurag@Gurome » Sat Jan 28, 2012 2:26 am
melon wrote:A bag contains 3 white balls, 3 black balls & 2 red balls. One by one three balls are drawn out without replacement. What is the probability that the third ball is red?
The probability of picking a red ball is 2/8 = 1/4 = 0.25, and probability of picking the red will not change for the successive draws, whether it is the 2nd, 3rd, or the 4th draw.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by mankey » Sat Jan 28, 2012 8:48 am
Dear Anurag

This question is without replacement, your approach would have been right if it were with replacement since it is possible that both the red balls get picked in the first two positions and there is no red ball left for the final position.

Please clarify.

Regards.
Join the discussion

by LalaB » Sat Jan 28, 2012 8:52 am
mankey wrote:Dear Anurag

This question is without replacement, your approach would have been right if it were with replacement since it is possible that both the red balls get picked in the first two positions and there is no red ball left for the final position.

Please clarify.

Regards.
actually, the answer will be the same with or without replacement :)

just think - u have some balls in front of u. u are not said the probability of first 2 balls. if u were said, then u would have been right, since first 2 links to the 3d one. but here u have some balls, and nothing is said, so all these balls have the equal right to be the 3d one

moreover, if the q. is what is the probability that the 1st ball is red, u again get the same answer :)
Join the discussion